2 (x + 1)^2/3 - 3 (x+1)^1/3 - 2 = 0
(2 answers)

Answer by MissionPossible(18)

2 (x + 1)^2/3 - 3 (x+1)^1/3 - 2 = 0
suppose (x+1)^1/3 = A
2A^2 - 3A - 2 = 0
(2A+1)(A-2) = 0
A = -1/2, 2
(x+1)^1/3 = -1/2
x+1 = (-1/2)^3
x+1 = -1/8
x = -9/8
(x+1)^1/3 = 2
x+1 = 2^3
x+1 = 8
x = 7
then x = -9/8 , 7 Ans.